OpenStudy (anonymous):

An equation that goes through point (1.4) and is parallel to y=-3/4x+2?

OpenStudy (anonymous):

suppose to be *-3/4x. yes i know the slope. I got a y= -3/4x +17/12. The book says y=-3/4x +19/4

OpenStudy (chihiroasleaf):

so, the equation on the question is \[y=-\frac{ 3 }{ 4 }x + 2\] ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i got 5/4 for y and -8/3 for x.... i don't see any other way to get another answer.

OpenStudy (chihiroasleaf):

ahh.., the lines are parallel... so, the slope is -3/4 since they are parallel, they have the same slope then just use the formula \[y-y _{1}=m \left( x-x _{1} \right)\] \[(y _{1},x _{1})\] is the point

OpenStudy (anonymous):

i did.... and i got a different answer from the book.

OpenStudy (anonymous):

i solved for the new x and y, put them into point slope form, then i tried to check if it's correct in normal slope form, but my answer is not the same as in the book

OpenStudy (chihiroasleaf):

\[y _{1} = 4 ; x _{1}=1\] so, y - 1 = (-3/4)(x - 1) y = (-3/4)x + (3/4) + 1 y = (-3/4)x + 7/4 right? *maybe the answer on the book is not correct :p

OpenStudy (anonymous):

i think my x is wrong since i plugged in what i got for x and it isn't 4

OpenStudy (anonymous):

no... why did you subtract y and 1? the y value given is 4...

OpenStudy (anonymous):

nvm... x value is -8/3 and the y vaLaue is 5/4... however i still don't get 19/4

OpenStudy (chihiroasleaf):

ahh..., I did a mistake..., sorry.. Yes.., it should be 4..., sorry..

OpenStudy (chihiroasleaf):

y - 4 = (-3/4)(x - 1) y = (-3/4)x + (3/4) + 4 y = (-3/4)x + 19/4

OpenStudy (anonymous):

oh ok thanks. i over complicated the problem

OpenStudy (chihiroasleaf):

you're welcome... :)